# Learning Passive Continuous-Time Dynamics with Multistep Port-Hamiltonian Gaussian Processes

### v1

@@ I Introduction
Learning physically consistent continuous-time dynamics is a core requirement for modeling and control, but noise and irregular sampling complicate the identification. Gaussian processes (GPs) provide a principled Bayesian framework for uncertainty quantification in system identification [1, 2], yet standard GP approaches typically ignore important structure such as energy conservation and passivity that are intrinsic to many mechanical and electrical systems. Recent work has begun to combine GP learning with port-Hamiltonian system (PHS) constraints, injecting physics-informed inductive biases via tailored kernels [3, 4]. These approaches directly enforce the port-Hamiltonian [5] structure in the learned dynamical model. However, these methods do not fully address the practical obstacles of measurement noise and irregular sampling in trajectories.
@@ I-A Related Works
Gaussian processes (GPs) offer nonparametric priors with closed-form posteriors under linear functionals and have been used effectively for continuous-time vector-field recovery [11]. In many pipelines, derivative preprocessing and regular sampling are pragmatic choices that work well [3, 12, 13]. At the same time, when samples are irregular and end-to-end uncertainty quantification is essential, it is appealing to bring discretization inside the statistical model rather than treating it as a separate preprocessing stage.
The vLMM schemes such as AB/AM/BDF [14] translate flows into linear constraints in stacked vector field evaluations. Recent work shows that these constraints can be projected into GP posteriors for exact inference, even with variable step size [7]. This line of work substantially advances GP-based identification under irregular sampling. A key contribution of our work is to integrate physics-informed PHS priors [3] into the vLMM-based GP [7] in order to incorporate the noise and sampling irregularity into the posteriors of the Hamiltonian surface, enabling geometric uncertainty quantification with meaningful physical interpretations.
@@ I-B Notations
The notation $\mathbb{R}$ denotes the real number line. Vectors in $\mathbb{R}^{n}$ are column vectors. The kronecker product is denoted as $\otimes$. The notation $\rm vec$} denote the stacking operator that turns $X\in\mathbb{N}^{m\times n}$ into ${\rm vec}(X)\in\mathbb{R}^{mn}$. $A]_{ij}$ denotes the $i,j$ entry of a matrix $A$. We write ${\rm det}(\cdot)$ for the determinant of a matrix. Gradient of scalar function $H$ with respect to $x$ is denoted as $\nabla_{x}H$. We write $\mathbb{E}[\cdot]$ for expectation. The Gaussian process distribution is denoted as $\mathcal{GP}(\cdot,\cdot)$.
@@ II-A Port-Hamiltonian Systems
in which state ${x}(t)\in\mathbb{R}^{n}$ and the I/O ports ${u}(t),{y}_{\rm out}(t)\in\mathbb{R}^{m}$ evolve accordingly with time $t\in\mathbb{R}_{\geq 0}$. The Hamiltonian $H:\mathbb{R}^{n}\to\mathbb{R}$ is a smooth function that represents the total energy stored in the system. The interconnection matrix $J:\mathbb{R}\to\mathbb{R}^{n\times n}$ is skew-symmetric. The dissipative matrix $R:\mathbb{R}\to\mathbb{R}^{n\times n}$ is a positive semi-definite such that $R({x})=R^{\top}({x})\succeq 0$. The port mapping matrix $G:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ defines how an external input/output ports u{u}, ${y}_{\rm out}$ are coupled to the energy storage dynamics.
@@ II-B Gaussian Process Regression (GPR)
A Gaussian process (GP) is a Bayesian prior over functions $f:\mathbb{R}^{d}\to\mathbb{R}$, where $f\sim\mathcal{GP}(m(\cdot),k(\cdot,\cdot))$, such that any finite collection of function values is jointly Gaussian [1]. The GP distribution, $\mathcal{GP}(m(\cdot),k(\cdot,\cdot))$, is fully specified by the mean and kernel function:
@@
$\begin{split}\mu_{*}&=m(X_{*})+K_{*X}\bigl(K_{XX}+\sigma_{y}^{2}I\bigr)^{-1}({y}-m(X)),\\ \Sigma_{*}&=K-K_{*X}\bigl(K_{XX}+\sigma_{y}^{2}I\bigr)^{-1}K_{X*},\end{split}$ (4)
@@ II-B1 Port–Hamiltonian System Kernel
A zero‐mean GP prior is propose in [3] to place on the unknown Hamiltonian:
@@
where $J_{R}({x})=J({x})-R({x})$ encodes the skew‐symmetric interconnection and dissipation matrices, and $\sigma_{f}^{2}\in\mathbb{R}_{>0}$ encode the signal noise. The above setup yields a GP prior:
@@ III-B Multistep Port-Hamiltonian System Kernel
To perform exact Gaussian process inference of a Port–Hamiltonian vector field with irregularly-sampled trajectory data points, we combine the physics‐informed PHS prior from [3] with the variable step size multistep integrator  [7].
@@ III-B1 Multivariate Variable Step Size Multistep Projection
Given noisy state observations $\{\tilde{{x}}_{k}\}_{k=1}^{K}$ at times $\{t_{k}\}_{k=1}^{K}$, an order-$(p)$ multistep integrator, e.g. Adams–Bashforth [14, Ch. 3.5], defines coefficient matrices, $A,B\in\mathbb{R}^{(K-M)\times K},$ such that the dynamic constraints of a control-affine system:
$\dot{x}(t)=f(x(t))+g(x(t))u(t),$
with $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$, ${g}:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ and $n=1$, can be enforced through:
@@
with the noiseless snapshot $X=[x_{1},\dots,x_{K}]^{\top}$ and $U=[{u}_{1},\dots,{u}_{K}]^{\top}$. The functions, $\mathbf{f}(X)=[f(x_{1}),\dots,f(x_{K})]^{\top}$ and $\mathbf{g}(X)U=[g(x_{1}){u}_{1},\dots,g(x_{K}){u}_{K}]^{\top},$ stacks the values of $f(x_{k})$ and $g(x_{k})u_{k}$, for $k=1\dots K$, accordingly. Notice that the first‐order forward Euler method with step size $h_{k}=t_{k+1}-t_{k}$ for zero-input systems is a special case of (12).
@@
When ${x}_{k}\in\mathbb{R}^{n}$ for $n\in\mathbb{N}$, the dynamic constraint encoded in (12) can be conveniently generalized to:
$A_{I}\coloneq A\otimes\mathsf{I}_{n},\ \$$B_{I}\coloneq B\otimes\mathsf{I}_{n},$ (13)
where $A_{I},B_{I}\in\mathbb{R}^{(K-M)n\times Kn}$, such that:
$Y\coloneq A_{I}{\rm vec}{(\tilde{X})}=B_{I}{\rm vec}{(\mathbf{f}(X)+\mathbf{g}(X)U)}+\boldsymbol{\varepsilon}.$ (14)
@@
We now extend the exact GP inference framework proposed in [7] by injecting a physics-informed prior via the matrix-value PHS kernel and construct the multistep port-Hamiltonian system (MS-PHS) kernel:
$[K_{Y}]_{nm}$ $=\sum_{i=0}^{K-M}\sum_{j=0}^{K-M}[B_{I}]_{n,i}k_{\rm phs}({x}_{i},{x}_{j})[B_{I}]_{m,j},$ (15)
$[k_{Y}({x}_{*})]_{n}$ $=\sum_{j=0}^{K-M}[B_{I}]_{n,j}k_{\rm phs}({x}_{*},{{x}}_{j}),$ (16)
@@
MS-PHS inherits passivity and closure under power-conserving interconnection from GP-PHS [3, Prop. 1,2], a direct consequence of the closure of Gaussian processes under linear operators. Conditioning the GP prior on the multistep linear functionals yields a posterior GP with the same smoothness, so the gradients required for energy-rate and passivity calculations remain well-defined. Since the multistep projection acts componentwise and linearly on the port variables, it preserves energy additivity and the Port-Hamiltonian interconnection structure.
@@ III-B2 Exact Inference via Multistep PHS Kernel
$\Sigma_{{f}}$ $=k_{\rm phs}(x^{*},x^{*})-k_{Y}({x}_{*})^{\top}\bigl(K_{Y}+\sigma_{{x}}^{2}A_{I}A_{I}^{\top}\bigr)^{-1}k_{Y}({x}_{*}),$ (20)
@@
where $\ell>0$ is the number of data points, $\widehat{\Phi},\Phi_{\rho}\in\mathcal{W}$ are the estimated and target flow of (10). Combining (22) with the vLMM LTE (21) produces a two-term high-probability vector-field bound:
@@
Since Tikhonov regularization, e.g., GPR, has Lipschitz exponent $\mu=2$, we take $\beta=\max\{1,2\mu\}=4$. Under the assumption that $H$ lies in the base RKHS induced by $k_{\rm base}$, the baseline source condition $r=\frac{1}{2}$, i.e., $\Phi_{\rho}\in\mathrm{Range}({\mathsf{T}}^{\frac{1}{2}}))$ makes [6, Thm. 2] yield the rate exponent $2r/(2r+\beta)=\frac{1}{5}$. If stronger smoothness justifies $r=2$, the exponent improves to $2r/(2r+\beta)=\frac{1}{2}$ [20, Sec 3.2]. ∎
@@ III-C Posterior Hamiltonian Surface
    $K_{ff}(X,X)=K_{Y}+\sigma_{{x}}^{2}A_{I}A_{I}^{\top}.$
@@
    $K_{Hf}(0,X)=\begin{bmatrix}(J-R)(x_{1}),\nabla_{x_{j}}k_{\rm base}(0,x_{1})\\ \vdots\\ (J-R)(x_{1}),\nabla_{x_{K}}k_{\rm base}(0,x_{K})\end{bmatrix}B_{I}^{\top},$ (25)
@@
$K_{*g}({x}_{*})=\bigl[k_{\rm base}({x}_{*},0),K_{Hf}({x}_{*},X)\bigr]\quad\in\mathbb{R}^{1+Kn}.$
In particular, $K_{Hf}({x}_{*},X)$ is constructed exactly as (25) but replacing the anchor with the test input ${x}_{*}$. Equations (27)–(28) then yield closed‐form expressions for the posterior mean and high‐probability variance of the Hamiltonian surface.
@@ IV-A Benchmarks
representing a undamped harmonic oscillator.
@@
To generate training and test data, each system is integrated over $t\in[0,20]$ using a classical fourth‐order Runge–Kutta solver with fixed time step $\Delta t=4\times 10^{-3}$. The resulting trajectories are then corrupted by additive state noise drawn from $\mathcal{N}(0,\sigma_{x}^{2}I)$ and irregularly subsampled at time points $t_{k}\in[0,20]$. For the irregularly subsampling scheme, we start from an $N$-point uniform grid $\{\tau_{k}\}_{k=1}^{N}$ on $t_{0},t_{1}]$ and add i.i.d. Gaussian “jitter” $\epsilon_{k}\sim\mathcal{N}\left(0,\sigma_{j}^{2}\right)$ with $\sigma_{j}=0.05$. The irregular timestamps are $t_{k}=\mathrm{clip}\left(\tau_{k}+\epsilon_{k},t_{0},t_{1}\right),$ after which they are sorted to enforce monotonicity. This produces near-uniform sampling with small, zero-mean perturbations, allowing mild clustering/gaps without changing the average sampling rate. Furthermore, we apply zero input ($u(t)\equiv 0$) on the Van der Pol oscillator and a nominal sinusoidal $\cos(\omega t)$ on the duffing oscillator and mass-spring system. We compare three inference schemes:
@@
    MS-PHS + Adams–Bashforth (MS-PHS-ab): our proposed multistep Port–Hamiltonian GP with Adams–Bashforth orders 1-3 [14, Ch. 3];
@@
    MS-ODE + Adams–Bashforth (MS-ODE-ab): the same multistep GP exact inference without a PHS prior with Adams–Bashforth orders 1-3, similar to the method proposed in [7];
@@
    GP-PHS + Savitzky–Golay (GP-PHS-savgol): a baseline that pre-processes the noisy data with a Savitzky–Golay filter [26] before fitting a PHS kernel [3]. We use Savitzky–Golay order 3 (savgol-3) in this paper.
  4. 4.
    GP-PHS + LOESS (GP-PHS-loess): a baseline that pre-processes the noisy data with locally estimated scatterplot smoothing (LOESS), a.k.a. locally weighted regression (LOWESS) [27, 28], before fitting a PHS kernel [3]. We use locally quadratic fitting (loess-2) in our simulations.

### v2

@@ I Introduction
Learning physically consistent continuous-time dynamics is a core requirement for modeling and control, but noise and irregular sampling complicate identification. Gaussian processes (GPs) provide a principled Bayesian framework for uncertainty quantification in system identification [1, 2], yet standard GP approaches typically ignore important structure such as energy conservation and passivity that are intrinsic to many mechanical and electrical systems. Recent work has begun to combine GP learning with port-Hamiltonian system (PHS) constraints, injecting physics-informed inductive biases via tailored kernels [3, 4]. These approaches directly enforce the port-Hamiltonian [5] structure in the learned dynamical model. However, these methods do not fully address the practical obstacles of measurement noise and irregular sampling in trajectories.
@@ I-A Related Works
Gaussian processes offer nonparametric priors with closed-form posteriors under linear functionals and have been used effectively for continuous-time vector-field recovery [11]. In many pipelines, derivative preprocessing and regular sampling are pragmatic choices that work well [3, 12, 13]. At the same time, when samples are irregular and end-to-end uncertainty quantification is essential, it is appealing to bring discretization inside the statistical model rather than treating it as a separate preprocessing stage.
The vLMM schemes such as AB/AM/BDF [14] translate flows into linear constraints in stacked vector field evaluations. Recent work shows that these constraints can be projected into GP posteriors for exact inference, even with variable step size [7]. This line of work substantially advances GP-based identification under irregular sampling. A key contribution of our work is to integrate physics-informed PHS priors [3] into the vLMM-based GP [7] to propagate noise and sampling irregularity into the Hamiltonian posterior. Our proposed GP framework enables the quantification of geometric uncertainty with meaningful physical interpretations.
@@ I-B Notations
The notation $\mathbb{R}$ denotes the real number line. Vectors in $\mathbb{R}^{n}$ are column vectors. The Kronecker product is denoted as $\otimes$. The matrix $I_{n}$ denotes the $n\times n$ identity. The operator $\rm vec$} denotes column-wise vectorization, mapping $X\in\mathbb{N}^{m\times n}$ into ${\rm vec}(X)\in\mathbb{R}^{mn}$. $A]_{ij}$ denotes the $i,j$ entry of a matrix $A$. We write ${\rm det}(\cdot)$ for the determinant of a matrix. Gradient of scalar function $H$ with respect to $x$ is denoted as $\nabla_{x}H$. We write $\mathbb{E}[\cdot]$ for expectation. The GP distribution is denoted as $\mathcal{GP}(\cdot,\cdot)$.
@@ II-A Port-Hamiltonian Systems
in which state ${x}(t)\in\mathbb{R}^{n}$ and the I/O ports ${u}(t),{y}_{\rm out}(t)\in\mathbb{R}^{m}$ evolve accordingly with time $t\in\mathbb{R}_{\geq 0}$. The Hamiltonian $H:\mathbb{R}^{n}\to\mathbb{R}$ is a smooth function that represents the total energy stored in the system. The interconnection matrix $J:\mathbb{R}^{n}\to\mathbb{R}^{n\times n}$ is skew-symmetric. The dissipative matrix $R:\mathbb{R}^{n}\to\mathbb{R}^{n\times n}$ is a positive semi-definite such that $R({x})=R^{\top}({x})\succeq 0$. The port mapping matrix $G:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ defines how the external input/output ports u{u}, ${y}_{\rm out}$ are coupled to the energy storage dynamics.
@@ II-B Gaussian Process Regression (GPR)
A GP is a Bayesian prior over functions $f:\mathbb{R}^{d}\to\mathbb{R}$, where $f\sim\mathcal{GP}(m(\cdot),k(\cdot,\cdot))$, such that any finite collection of function values is jointly Gaussian [1]. The GP distribution, $\mathcal{GP}(m(\cdot),k(\cdot,\cdot))$, is fully specified by the mean and kernel function:
@@
$\begin{split}\mu_{*}&=m(X_{*})+K_{*X}\bigl(K_{XX}+\sigma_{y}^{2}I\bigr)^{-1}({y}-m(X)),\\ \Sigma_{*}&=K_{**}-K_{*X}\bigl(K_{XX}+\sigma_{y}^{2}I\bigr)^{-1}K_{X*},\end{split}$ (4)
@@ II-B1 Port–Hamiltonian System Kernel
A zero‐mean GP prior is proposed in [3] to place on the unknown Hamiltonian:
@@
where $J_{R}({x})=J({x})-R({x})$ encodes the interconnection and dissipation matrices, and $\sigma_{f}^{2}\in\mathbb{R}_{>0}$ encodes the signal noise. The above setup yields a GP prior:
@@ III-A Problem Formulation
Under the above setup, our goal is to learn a probabilistic, physics-consistent model of the zero-input drift $f(x)=[J(x)-R(x)]\nabla_{x}H(x)$ and of the Hamiltonian surface $H(x)$ that: (i) combines vLMM constraints and PHS structure in a single GP likelihood; (ii) derives closed-form posterior mean and covariance for both the vector field and the Hamiltonian at arbitrary states, with $H$ identified up to an additive constant fixed by an anchor $H(x_{0})=H_{0}$; and (iii) admits a finite-sample, high-probability error bound for the vector-field estimate that separates statistical fit from numerical discretization, exposing the roles of data points $\ell=K-M$ and multistep accuracy. Kernel and PHS hyperparameters are selected by minimizing (5).
@@ III-B Multistep Port-Hamiltonian System Kernel
To perform exact GP inference of a Port–Hamiltonian vector field with irregularly-sampled trajectory data points, we combine the physics‐informed PHS prior from [3] with the variable step size multistep integrator  [7].
@@ III-B1 Multivariate Variable Step Size Multistep Projection
With noisy state observations $\{\tilde{{x}}_{k}\}_{k=1}^{K}$ at time steps $\{t_{k}\}_{k=1}^{K}$, we define an order-$(p)$ vLMM [14, Ch. 3.5] with the coefficient matrices, $A,B\in\mathbb{R}^{(K-M)\times K}$, parameterized by the step grid $\{h_{k}\}_{k=1}^{K-1}$, where $h_{k}=t_{k+1}-t_{k}$. Let the control-affine system be:
$\dot{x}(t)=f(x(t))+g(x(t))u(t),$ (12)
with $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$, ${g}:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$. In the scalar case where $n=1$, dynamic constraints in (12) can be enforced through:
@@
with the noiseless states $X=[x_{1}^{\top},\dots,x_{K}^{\top}]^{\top}$ and inputs $U=[{u}_{1}^{\top},\dots,{u}_{K}^{\top}]^{\top}$. The functions, $\mathbf{f}(X)=[f(x_{1})^{\top},\dots,f(x_{K})^{\top}]^{\top}$ and $\mathbf{g}(X)U=[(g(x_{1}){u}_{1})^{\top},\dots,(g(x_{K}){u}_{K})^{\top}]^{\top},$ stacks the values of $f(x_{k})$ and $g(x_{k})u_{k}$, for $k=1\dots K$, accordingly. Notice that the forward Euler method with step sizes $\{h_{k}\}$ for zero-input systems is a special case of (13).
@@
In the multivariate case where $x\in\mathbb{R}^{n}$ for $n>1$, the dynamic constraint encoded in (13) can be conveniently generalized by augmenting the coefficient matrices:
$A_{I}\coloneq A\otimes I_{n},\ \$$B_{I}\coloneq B\otimes I_{n},$ (14)
where $A_{I},B_{I}\in\mathbb{R}^{(K-M)n\times Kn}$. The label vector $Y$ for GPR is constructed and related to the feature vector $X$ via:
$Y\coloneq A_{I}\tilde{X}=B_{I}{(\mathbf{f}(X)+\mathbf{g}(X)U)}+\boldsymbol{\varepsilon}.$ (15)
@@
We now extend the exact GP inference framework proposed in [7] by injecting a physics-informed prior via the matrix-valued PHS kernel and construct the multistep port-Hamiltonian system (MS-PHS) kernel:
$[K_{Y}]_{nm}$ $=\sum_{i=0}^{K}\sum_{j=0}^{K}[B_{I}]_{n,i}k_{\rm phs}({x}_{i},{x}_{j})[B_{I}]_{m,j},$ (16)
$[k_{Y}({x}_{*})]_{n}$ $=\sum_{j=0}^{K}[B_{I}]_{n,j}k_{\rm phs}({x}_{*},{{x}}_{j}),$ (17)
@@
MS-PHS inherits passivity and closure under power-conserving interconnection from GP-PHS [3, Prop. 1,2], a direct consequence of the closure of GPs under linear operators. Conditioning the GP prior on the multistep linear functionals yields a posterior GP with the same smoothness, so the gradients required for energy-rate and passivity calculations remain well-defined. Since the multistep projection acts componentwise and linearly on the port variables, it preserves energy additivity and the Port-Hamiltonian interconnection structure.
@@ III-B2 Exact Inference via Multistep PHS Kernel
$\Sigma_{{f}}$ $=k_{\rm phs}(x_{*},x_{*})-k_{Y}({x}_{*})^{\top}\bigl(K_{Y}+\sigma_{{x}}^{2}A_{I}A_{I}^{\top}\bigr)^{-1}k_{Y}({x}_{*}),$ (21)
@@
where $\ell>0$ is the number of labeled windows $\ell\coloneqq K-M$, $\widehat{\Phi},\Phi_{\rho}\in\mathcal{W}$ are the estimated and target flow of (10). Combining (23) with the vLMM LTE (22) produces a two-term high-probability vector-field bound:
@@
Since Tikhonov regularization, e.g., GPR, has Lipschitz exponent $\mu=2$, we take $\beta=\max\{1,2\mu\}=4$. Under the assumption that $H$ lies in the base RKHS induced by $k_{\rm base}$, the baseline source condition $r=\frac{1}{2}$, i.e., $\Phi_{\rho}\in\mathrm{Range}({\mathsf{T}}^{\frac{1}{2}})$ makes [6, Thm. 2] yield the rate exponent $2r/(2r+\beta)=\frac{1}{5}$. If stronger smoothness justifies $r=2$, the exponent improves to $2r/(2r+\beta)=\frac{1}{2}$ [20, Sec. 3.2]. ∎
@@ III-C Posterior Hamiltonian Surface
    $K_{ff}(X,X)=K_{Y}+\sigma_{{x}}^{2}A_{I}A_{I}^{\top},$
    which matches the training covariance in (20) and (21).
@@
    $K_{Hf}(0,X)=\begin{bmatrix}J_{R}(x_{1})\nabla_{x_{1}}k_{\rm base}(0,x_{1})\\ \vdots\\ J_{R}(x_{K})\nabla_{x_{K}}k_{\rm base}(0,x_{K})\end{bmatrix}B_{I}^{\top},$ (26)
@@
$K_{*g}({x}_{*})=\bigl[k_{\rm base}({x}_{*},0),K_{Hf}({x}_{*},X)\bigr]\quad\in\mathbb{R}^{1\times(1+(K-M)n)}.$
In particular, $K_{Hf}({x}_{*},X)$ is constructed exactly as (26) but replacing the anchor with the test input ${x}_{*}$. Equations (28)–(29) then yield closed‐form expressions for the posterior mean and variance of the Hamiltonian surface.
@@ IV-A Benchmarks
representing an undamped harmonic oscillator.
@@
To generate training and test data, each system is integrated over $t\in[0,20]$ using a classical fourth‐order Runge–Kutta solver with fixed time step $\Delta t=4\times 10^{-3}$. The resulting trajectories are then corrupted by additive state noise drawn from $\mathcal{N}(0,\sigma_{x}^{2}I)$ and irregularly subsampled at time points $t_{k}\in[0,20]$. For the irregularly subsampling scheme, we start from an $N$-points uniform grid $\{\tau_{k}\}_{k=1}^{N}$ on $[0,20]$ with $N=100$ and add i.i.d. Gaussian “jitter” $\epsilon_{k}\sim\mathcal{N}\left(0,\sigma_{j}^{2}\right)$ with $\sigma_{j}=0.05$. The irregular timestamps are $t_{k}=\mathrm{clip}\left(\tau_{k}+\epsilon_{k},t_{0},t_{1}\right),$ after which they are sorted to enforce monotonicity. This produces near-uniform sampling with small, zero-mean perturbations, allowing mild clustering/gaps without changing the average sampling rate. Furthermore, we apply zero input ($u(t)\equiv 0$) on the Van der Pol oscillator and a nominal sinusoidal $\cos(\omega t)$ on the Duffing oscillator and mass-spring system. We compare three inference schemes:
@@
    MS-PHS + Adams-Bashforth (MS-PHS-ab): our proposed multistep Port–Hamiltonian GP with Adams–Bashforth orders 1-3 [14, Ch. 3];
@@
    MS-ODE + Adams-Bashforth (MS-ODE-ab): the same multistep GP exact inference without a PHS prior with Adams–Bashforth orders 1-3, similar to the method proposed in [7];
@@
    GP-PHS + LOESS/Savitzky-Golay (GP-PHS-loess/savgol): a baseline that pre-processes the noisy data with a locally estimated scatterplot smoothing (LOESS), a.k.a. locally weighted regression (LOWESS) [26, 27] before fitting a PHS kernel [3]. The Savitzky-Golay filter [28] can be seen as a regular grid variant of LOESS. We use locally quadratic fitting (loess-2) and Savitzky-Golay order 3 (savgol-3) in our simulations.